OpenStudy (anonymous):

Which of theses values is equivalent to the change in momentum of an object? (A)Impulse (B)acceleration (C)force (D)velocity

OpenStudy (souvik):

impulse =change in momentum...

OpenStudy (anonymous):

it would be impulse if time is taken into consideration

OpenStudy (souvik):

F=ma [F=force ,m=mass, a= acceleration] F=m*(v-u)/t F*t=mv-mu impulse=change in momentum...

OpenStudy (anonymous):

yes thank you i got it right :) can you guys help me with the rest of the questions?

OpenStudy (souvik):

sure..